# DREAMS: Preserving both Local and Global Structure in Dimensionality Reduction

### v1

@@ Abstract
Dimensionality reduction techniques are widely used for visualizing high-dimensional data in two dimensions. Existing methods are typically designed to preserve either local (e.g. $t$-SNE, UMAP) or global (e.g. MDS, PCA) structure of the data, but none of the established methods can represent both aspects well. In this paper, we present DREAMS (Dimensionality Reduction Enhanced Across Multiple Scales), a method that combines the local structure preservation of $t$-SNE with the global structure preservation of PCA via a simple regularization term. Our approach generates a spectrum of embeddings between the locally well-structured $t$-SNE embedding and the globally well-structured PCA embedding, efficiently balancing both local and global structure preservation. We benchmark DREAMS across seven real-world datasets, including five from single-cell transcriptomics and one from population genetics, showcasing qualitatively and quantitatively its superior ability to preserve structure across multiple scales compared to previous approaches.
@@ 1 Introduction
Real-world data often exhibits high-dimensional and complex structures, making their effective and interpretable visualization a crucial step in exploratory data analysis. Dimensionality reduction methods serve this purpose by projecting high-dimensional data into more interpretable low-dimensional representations while preserving meaningful structures. Among dimensionality reduction methods, principal component analysis (PCA, Hotelling 1933) and $t$-distributed stochastic neighbor embedding ($t$-SNE, van der Maaten and Hinton 2008) have emerged as two widely used methods, each with distinct strengths. PCA excels at capturing global structures by projecting data onto a lower-dimensional subspace in directions that maximize variance, providing a broad overview of the dataset’s structure. In contrast, $t$-SNE is a neighbor-embedding method with the objective to map data points that are nearby in high-dimensional space close to one another in the low-dimensional embedding. This focus on neighbor preservation makes $t$-SNE particularly effective in preserving local structures. When applied individually, both methods suffer from limitations: PCA often overlooks fine local relationships, while $t$-SNE distorts global structures in favor of local neighborhood preservation.
@@
However, in many real-world datasets, e.g. single-cell transcriptomic datasets, both local and global structure is meaningful. Local relationships can reveal microscopic data patterns such as small clusters and cell types. For example, the Tasic et al. dataset features 122 fine clusters that represent cell types, which are clearly separated in the $t$-SNE plot (Figure 1 right). Conversely, the global structure reflects macroscopic data patterns such as separations between broad cell classes (Tasic et al., 2018) or developmental trajectories (Kanton et al., 2019) across an entire cell population. In the Tasic et al. dataset, this level of structure separates individual cells into non-neurons, excitatory, and inhibitory neurons. This global structure is not evident in the $t$-SNE plot, but is prominent in the PCA visualization, which in turn fails to separate the finer clusters (Figure 1 left). Since both local and global structure carry essential and often complementary information, neglecting either structure scale can result in incomplete or biased interpretations, highlighting the importance of preserving both scales simultaneously.
@@
Our proposed method DREAMS combines the interpretability and global structure preservation of PCA with the local sensitivity of $t$-SNE in a simple, yet effective way. On the Tasic et al. data, it maintains PCA’s global arrangement by non-neurons, inhibitory, and excitatory neurons, while also revealing the finer cluster structure within each broad group, similar to $t$-SNE (Figure 1 middle). DREAMS integrates global structure preservation into the $t$-SNE objective by a PCA-based regularization term applied throughout the entire optimization process. Varying the regularization strength allows DREAMS to transition from the locally well-structured $t$-SNE embedding to the globally well-structured PCA embedding. Along this spectrum it trades off local and global structure more favorably than its competitors resulting in improved qualitative and quantitative structure preservation, particularly on datasets with hierarchical organization that exhibit both prominent local and global patterns.
@@
Figure 1: PCA, DREAMS, and $t$-SNE embeddings of the Tasic et al. dataset (Tasic et al., 2018) illustrate how DREAMS preserves the global organization seen in the PCA embedding — such as the separation of non-neurons, inhibitory, and excitatory neurons — while also capturing the local cell-type structure that is present in the $t$-SNE embedding.
@@
    present a spectrum of visualizations from $t$-SNE to PCA with almost no local/global structure trade-off;
@@
    perform a benchmark of ten algorithms on seven real-world datasets;
@@
Our code to reproduce the results of this paper is available at https://github.com/berenslab/dreams-experiments. We used a modified openTSNE implementation available at https://github.com/berenslab/DREAMS and a modified CNE implementation available at https://github.com/berenslab/DREAMS-CNE.
@@ 2 Related Work
The most prominent and well-established visualization methods either excel at preserving local structure, e.g., neighbor-embedding methods, such as $t$-SNE (van der Maaten and Hinton, 2008) and UMAP (McInnes et al., 2018), or global structure such as PCA (Pearson, 1901; Hotelling, 1933) and MDS (Kruskal, 1964), but not both simultaneously. Therefore, many recent efforts have aimed at producing visualizations with faithful local and global structure.
Most of them attempt to improve the global structure preservation of neighbor embeddings. One way to make neighbor embeddings more sensitive to the global structure of a dataset is to widen the set of points that are deemed similar beyond a small set of nearest neighbors. For instance, increasing the perplexity parameter in $t$-SNE effectively increases the number of considered nearest neighbors but consequently leads to increased runtime (Kobak and Berens, 2019; Lee et al., 2015; De Bodt et al., 2020). Similarly, several modifications of UMAP, e.g. PaCMAP (Wang et al., 2021) and TriMAP (Amid and Warmuth, 2019), do not only use nearest neighbors, but also consider more distant points in their optimization. In addition to attraction between nearest neighbors, PaCMAP employs weak attraction on mid-near points, while TriMap tries to also preserve the order of similarities in random triplets. Another triplet-based method, ivis (Szubert et al., 2019), strives to balance local and global structure with a parametric encoder and a margin loss on triplets of embedding distances.
A different strategy to improve the global structure in neighbor embeddings is to initialize them with a global embedding, which can improve the global structure of the final embedding, despite no further steps for preserving global structure during optimization (Kobak and Linderman, 2021; Wang et al., 2021).
@@
More similar to our approach are non-parametric hybrid methods that directly optimize the embedding positions. Local-to-Global Structures (Miller et al., 2023), a method for generic graph drawing, applies MDS-like stress minimization to shortest path graph distances among pairs of points that are strongly connected, while repelling non-neighboring pairs. The SQuadMDS-hybrid (Lambert et al., 2022) interpolates between the objectives of $t$-SNE and MDS. For MDS it uses the stochastic quartet framework of Lambert et al. (2022). To make both objectives more compatible, SQuadMDS-hybrid normalizes their gradients before blending them together. The hybrid method StarMAP (Watanabe et al., 2025) changes the attractive force in the UMAP objective. Instead of pulling only nearest neighbors together, it also pulls points towards the PCA coordinates of their $k$-means cluster centroid. This combination of a neighbor embedding with PCA is similar to our DREAMS. But instead of UMAP, we use $t$-SNE due to its better local structure preservation. Moreover, we propose a simpler objective that avoids StarMAP’s clustering step, leaves the neighbor-embedding objective intact, and simply pulls each embedding point towards its own PCA position. We found that this simpler approach leads to a better local-global trade-off. Additionally, setting the regularization strength in DREAMS to 0 or to 1 allows to recover standard $t$-SNE and PCA, while the StarMAP framework cannot fully recover PCA.
@@ 3 Background
In dimensionality reduction, we aim to represent a high-dimensional dataset $X=[x_{1},\ldots,x_{n}]^{T}\in\mathbb{R}^{n\times m}$ with $n$ observations in an $m$-dimensional space by a lower-dimensional embedding $Y=[y_{1},\ldots,y_{n}]^{T}\in\mathbb{R}^{n\times d}$, where $d\ll m$. The objective is to construct $Y$ such that meaningful relationships among the observations are preserved in the lower-dimensional space. In this section, we outline the two dimensionality reduction methods that DREAMS builds upon — PCA and $t$-SNE — and show how they approach this goal from complementary perspectives.
@@ 3.1 Principal component analysis (PCA)
Principal component analysis (PCA; Pearson 1901; Hotelling 1933) is a linear transformation that projects the data onto a new coordinate system aligned with the directions of maximum variance. PCA seeks a linear, orthogonal mapping $W\in\mathbb{R}^{m\times d}$ that projects the data into a lower-dimensional space $Y=XW$, where the directions in $W$ capture the maximal variance across the entire dataset $X$. This ensures the preservation of the macroscopic, global structure since distances along directions with high variance are preserved in the projection (Huang et al., 2022). However, because orthogonal projections can map distant points to similar locations, PCA performs poorly at preserving local structures (Huang et al., 2022; Wang et al., 2023).
@@ 3.2 $t$-distributed stochastic neighbor embedding ($t$-SNE)
$t$-distributed stochastic neighbor embedding ($t$-SNE; van der Maaten and Hinton 2008) is a widely used neighbor embedding method, that is especially effective at preserving local similarities (Espadoto et al., 2019). By transforming Euclidean distances between points into pairwise similarity probabilities, $t$-SNE constructs a probability distribution, $P$, based on the high-dimensional observations $X$, and a probability distribution, $Q$, based on the low-dimensional embeddings $Y$. The distribution $P=\{p_{ij}\}^{n}_{i,j=1}$ encodes the nearest-neighbor structure in the high-dimensional space via
$p_{ij}=\frac{p_{j|i}+p_{i|j}}{2n},\textnormal{ where }p_{j|i}=\frac{\exp\left(-\|x_{i}-x_{j}\|^{2}/(2\sigma^{2}_{i})\right)}{\sum_{k\neq i}\exp\left(-\|x_{i}-x_{k}\|^{2}/(2\sigma^{2}_{i})\right)}\text{ if }i\neq j\text{ and }p_{i|i}=0.$
@@
The similarity probability distribution of the low-dimensional embedding points, $Q=\{q_{ij}\}^{n}_{i,j=1}$, is based on a Cauchy kernel:
$q_{ij}=\frac{\left(1+\|y_{i}-y_{j}\|^{2}\right)^{-1}}{\sum_{k\neq l}\left(1+\|y_{k}-y_{l}\|^{2}\right)^{-1}}\text{ if }i\neq j\text{ and }q_{ii}=0.$
@@
The similarity probabilities act as kernels centered around the data and embedding points, assigning high similarity values only to close neighbors while distant points have little impact on the loss function. This makes neighbors in the high-dimensional space remain neighbors in the low-dimensional embedding space, ensuring the preservation of local structure. In contrast, due to the weak influence of distant points, global structure can be misrepresented (Huang et al., 2022; Wang et al., 2023).
@@ 4.1 Regularizing with precomputed global embedding
In DREAMS, we first precompute the PCA positions $\widetilde{Y}\in\mathbb{R}^{n\times 2}$ as the reference embedding for global structure of the data. To combine the local structure preservation of $t$-SNE with the global structure preservation of PCA, we augment the $t$-SNE loss function by a regularization term that penalizes embedding points $Y$ deviating from their PCA positions $\widetilde{Y}$, yielding the loss
$\mathcal{L}(Y)=(1-\lambda)\mathcal{L}_{t\text{-SNE}}(Y)+\lambda\|Y-\alpha\widetilde{Y}\|_{2}^{2}.$ (2)
@@
$\alpha=\|Y\|_{2}\big{/}\|\widetilde{Y}\|_{2}.$ (3)
This scaling encourages the reference embedding to match the current scale of the $t$-SNE embedding during optimization, making the global and local objectives more compatible. Note that $\alpha$ was treated as a constant on each gradient descent iteration.
The first term of Equation 2 enforces the preservation of local structure in the embedding by minimizing the $t$-SNE loss (Equation 1), while the regularization term ensures the preservation of global structure by encouraging the embedding to resemble the (scaled) PCA embedding $\widetilde{Y}$. This setup allows for local adjustments by the $t$-SNE loss, while the quadratic penalty prevents large deviations that would distort the global layout. The regularization strength $\lambda$ controls the impact of the PCA embedding on the final embedding, thereby enabling a trade-off between local and global structure preservation. For $\lambda=0$, DREAMS produces a standard $t$-SNE embedding, while for $\lambda=1$, it outputs the standard PCA embedding. Empirically, we found $\lambda=0.1$ to be a suitable value for combining the individual strengths of global structure preservation of PCA with the local structure preservation of $t$-SNE without substantially compromising either aspect.
This framework is not limited to using the PCA embedding as $\widetilde{Y}$. Depending on the characteristics of the dataset, other methods with a good global structure preservation, such as MDS, can also be used as an effective choice for $\widetilde{Y}$. We will call the version of DREAMS that regularizes using the SQuadMDS embedding as DREAMS-MDS. By default, we use PCA instead of MDS because PCA is faster and is inherently interpretable due to its linearity.
Our implementation was based on the openTSNE library (Poličar et al., 2024) where we added the regularization term to the gradient descent optimization. We used default openTSNE hyperparameters and always initialized with the regularization embedding $\widetilde{Y}$. When using the PCA embedding as $\widetilde{Y}$, we normalized it by the standard deviation of PC1. When using the MDS embedding (DREAMS-MDS), we directly used the SQuadMDS output embedding, which performs some internal normalization and produces embeddings with standard deviations around 10 for all datasets.

### v2

@@ Abstract
Dimensionality reduction techniques are widely used for visualizing high-dimensional data in two dimensions. Existing methods are typically designed to preserve either local (e.g., $t$-SNE, UMAP) or global (e.g., MDS, PCA) structure of the data, but none of the established methods can represent both aspects well. In this paper, we present DREAMS (Dimensionality Reduction Enhanced Across Multiple Scales), a method that combines the local structure preservation of $t$-SNE with the global structure preservation of PCA via a simple regularization term. Our approach generates a spectrum of embeddings between the locally well-structured $t$-SNE embedding and the globally well-structured PCA embedding, efficiently balancing both local and global structure preservation. We benchmark DREAMS across eleven real-world datasets, showcasing qualitatively and quantitatively its superior ability to preserve structure across multiple scales compared to previous approaches.
Figure 1: PCA, DREAMS, and $t$-SNE embeddings of the Tasic et al. dataset illustrate how DREAMS preserves the global organization seen in the PCA embedding — such as the separation of non-neurons, inhibitory, and excitatory neurons — while also capturing the local cell-type structure that is present in the $t$-SNE embedding.
@@ 1 Introduction
Real-world data often exhibit complex structures, making their effective and interpretable visualization a crucial step in exploratory data analysis. Dimensionality reduction methods serve this purpose by projecting high-dimensional data into more interpretable low-dimensional representations while preserving meaningful structures (de Bodt et al., 2025). Among dimensionality reduction methods, principal component analysis (PCA; Hotelling, 1933) and $t$-distributed stochastic neighbor embedding ($t$-SNE; van der Maaten and Hinton, 2008) have emerged as two widely used methods. PCA excels at capturing global structures by projecting data onto a lower-dimensional subspace in directions that maximize variance, providing a broad overview of the dataset’s structure. In contrast, $t$-SNE is a neighbor-embedding method with the objective to map data points that are nearby in high-dimensional space close to one another in the low-dimensional embedding. Its focus on neighborhood preservation makes $t$-SNE particularly effective in preserving local structures. When applied individually, both methods suffer from limitations: PCA often overlooks fine local relationships, while $t$-SNE distorts global structures in favor of local neighborhood preservation.
@@
However, in many real-world datasets, e.g., single-cell transcriptomic datasets, both local and global structure is meaningful. Local relationships can reveal microscopic data patterns such as small clusters and cell types. For example, the Tasic et al. dataset features 122 fine clusters that represent cell types, which are clearly separated in the $t$-SNE plot (Figure 1 right). Conversely, the global structure reflects macroscopic data patterns such as separations between broad cell classes (Tasic et al., 2018) or developmental trajectories (Kanton et al., 2019) across an entire cell population. In the Tasic et al. dataset, this level of structure separates individual cells into non-neurons, excitatory, and inhibitory neurons. This global structure is not evident in the $t$-SNE plot, but is prominent in the PCA visualization, which in turn fails to separate the finer clusters (Figure 1 left). Since both local and global structure carry essential and often complementary information, neglecting either structure scale can result in incomplete or biased interpretations, highlighting the importance of preserving both scales simultaneously.
@@
Our proposed method DREAMS combines the interpretability and global structure preservation of PCA with the local sensitivity of $t$-SNE in a simple, yet effective way. On the Tasic et al. data, it maintains PCA’s global arrangement of non-neurons, inhibitory, and excitatory neurons, while also revealing the finer cluster structure within each broad group, similar to $t$-SNE (Figure 1 middle). DREAMS integrates global structure preservation into the $t$-SNE objective by a PCA-based regularization term applied throughout the entire optimization process. Varying the regularization strength allows DREAMS to transition from the locally well-structured $t$-SNE embedding to the globally well-structured PCA embedding. Along this spectrum it trades off local and global structure more favorably than its competitors, resulting in improved qualitative and quantitative structure preservation, particularly on datasets with hierarchical organization that exhibit both prominent local and global patterns.
@@
    present a spectrum of visualizations with state-of-the-art trade-off between local and global structure;
@@
    perform a benchmark of ten algorithms on eleven real-world datasets;
@@
Our code is available at https://github.com/berenslab/dreams-experiments/tree/tmlr. We used a modified openTSNE implementation available at https://github.com/berenslab/DREAMS/tree/tmlr and a modified CNE implementation available at https://github.com/berenslab/DREAMS-CNE/tree/tmlr.
@@ 2 Related Work
The most prominent and well-established visualization methods either excel at preserving local structure, e.g., neighbor-embedding methods $t$-SNE (van der Maaten and Hinton, 2008) and UMAP (McInnes et al., 2018), or global structure, e.g., PCA (Pearson, 1901; Hotelling, 1933) and MDS (Kruskal, 1964), but not both simultaneously. Therefore, many recent efforts have aimed at producing visualizations with faithful local and global structure.
Most of them attempt to improve the global structure preservation of neighbor embeddings. One way to make neighbor embeddings more sensitive to the global structure of a dataset is to widen the set of points that are deemed similar beyond a small set of nearest neighbors. For instance, increasing the perplexity parameter in $t$-SNE effectively increases the number of considered nearest neighbors but consequently leads to increased runtime (Kobak and Berens, 2019; Lee et al., 2015; De Bodt et al., 2020). Similarly, several modifications of UMAP, e.g., PaCMAP (Wang et al., 2021) and TriMAP (Amid and Warmuth, 2019), do not only use nearest neighbors, but also consider more distant points in their optimization. In addition to attraction between nearest neighbors, PaCMAP employs weak attraction on mid-near points, while TriMap tries to also preserve the order of similarities in random triplets. Another triplet-based method, ivis (Szubert et al., 2019), strives to balance local and global structure with a parametric encoder and a margin loss on triplets of embedding distances. Changing the similarity of input points more drastically, EmbedOR (Saidi et al., 2025) computes a neighbor embedding from a distance matrix based on the curvature, and thus geometry, of the $k$-nearest neighbor graph.
A different strategy to improve the global structure in neighbor embeddings is to initialize them with a global embedding, which can improve the global structure of the final embedding, despite no further steps for preserving global structure during optimization (Kobak and Linderman, 2021; Wang et al., 2021). UMATO (Jeon et al., 2025) first computes a skeletal layout of only the most densely connected points to capture the global layout and then adds the remaining points in a second optimization phase.
@@
More similar to our approach are non-parametric hybrid methods that directly optimize the embedding positions. Local-to-Global Structures (Miller et al., 2023), a method for generic graph drawing, applies MDS-like stress minimization to shortest path graph distances among pairs of points that are strongly connected, while repelling non-neighboring pairs. Cluster and Embed (Coda et al., 2025) embeds clusters separately, focusing on their local structure. These fixed cluster embeddings then get arranged into a full embedding by rigid transformations that minimize the overall stress like in MDS. The SQuadMDS-hybrid (Lambert et al., 2022) interpolates between the objectives of $t$-SNE and MDS. For MDS it uses the stochastic quartet framework of Lambert et al. (2022). To make both objectives more compatible, SQuadMDS-hybrid normalizes their gradients before blending them together. The hybrid method StarMAP (Watanabe et al., 2025) changes the attractive force in the UMAP objective. Instead of pulling only nearest neighbors together, it also pulls points towards the PCA coordinates of their $k$-means cluster centroid. This combination of a neighbor embedding with PCA is similar to our DREAMS. But instead of UMAP, we use $t$-SNE due to its better local structure preservation. Moreover, we propose a simpler objective that avoids StarMAP’s clustering step, leaves the neighbor-embedding objective intact, and simply pulls each embedding point towards its own PCA position. We found that this simpler approach leads to a better local-global trade-off. Additionally, setting the regularization strength in DREAMS to 0 or to 1 allows to recover standard $t$-SNE and PCA, while the StarMAP framework cannot fully recover PCA.
@@ 3 Background
In dimensionality reduction, we aim to represent a high-dimensional dataset $X=[x_{1},\ldots,x_{n}]^{\top}\in\mathbb{R}^{n\times m}$ with $n$ observations in an $m$-dimensional space by a lower-dimensional embedding $Y=[y_{1},\ldots,y_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $d\ll m$. The objective is to construct $Y$ such that meaningful relationships among the observations are preserved in the lower-dimensional space. In this section, we outline the two dimensionality reduction methods that DREAMS builds upon — PCA and $t$-SNE — and show how they approach this goal from complementary perspectives.
@@ 3.1 Principal component analysis (PCA)
Principal component analysis (PCA; Pearson, 1901; Hotelling, 1933) is a linear transformation that projects the data onto a new coordinate system aligned with the directions of maximum variance. PCA seeks a linear, orthogonal mapping $W\in\mathbb{R}^{m\times d}$ that projects the data into a lower-dimensional space $Y=XW$, where the directions in $W$ capture the maximal variance across the entire dataset $X$. This ensures the preservation of the macroscopic, global structure since distances along directions with high variance are preserved in the projection (Huang et al., 2022). However, because orthogonal projections can map distant points to similar locations, PCA performs poorly at preserving local structures (Huang et al., 2022; Wang et al., 2023).
@@ 3.2 $t$-distributed stochastic neighbor embedding ($t$-SNE)
$t$-distributed stochastic neighbor embedding ($t$-SNE; van der Maaten and Hinton, 2008) is a widely used neighbor-embedding method that is especially effective at preserving local similarities (Espadoto et al., 2019). By transforming Euclidean distances between points into pairwise similarity probabilities, $t$-SNE constructs a probability distribution $P$, based on the high-dimensional observations $X$, and a probability distribution $Q$, based on the low-dimensional embeddings $Y$. The distribution $P=\{p_{ij}\}^{n}_{i,j=1}$ encodes the nearest-neighbor structure in the high-dimensional space via
$p_{ij}=\frac{p_{j|i}+p_{i|j}}{2n},\textnormal{ where }p_{j|i}=\frac{\exp\left(-\|x_{i}-x_{j}\|^{2}/(2\sigma^{2}_{i})\right)}{\sum_{k\neq i}\exp\big(-\|x_{i}-x_{k}\|^{2}/(2\sigma^{2}_{i})\big)}\text{ if }i\neq j\text{ and }p_{i|i}=0.$
@@
The similarity probability distribution of the low-dimensional embedding points $Q=\{q_{ij}\}^{n}_{i,j=1}$ is based on the Cauchy kernel:
$q_{ij}=\frac{\left(1+\|y_{i}-y_{j}\|^{2}\right)^{-1}}{\sum_{k\neq l}\big(1+\|y_{k}-y_{l}\|^{2}\big)^{-1}}\text{ if }i\neq j\text{ and }q_{ii}=0.$
@@
The similarity probabilities act as kernels centered around the data and embedding points, assigning high similarity values only to close neighbors, while distant points have little impact on the loss function. This makes neighbors in the high-dimensional space remain neighbors in the low-dimensional embedding space, ensuring the preservation of local structure. In contrast, due to the weak influence of distant points, global structure can be misrepresented (Huang et al., 2022; Wang et al., 2023).
@@ 4.1 Regularizing with precomputed global embedding
In DREAMS, we first precompute the PCA positions $\widetilde{Y}\in\mathbb{R}^{n\times 2}$ as the reference embedding for the global structure of the data. To combine the local structure preservation of $t$-SNE with the global structure preservation of PCA, we augment the $t$-SNE loss function by a regularization term that penalizes embedding points $Y$ deviating from their PCA positions $\widetilde{Y}$, yielding the loss
$\mathcal{L}(Y)=(1-\lambda)\mathcal{L}_{t\text{-SNE}}(Y)+\frac{\lambda}{n}\|Y-\alpha\widetilde{Y}\|_{F}^{2}.$ (2)
@@
$\alpha=\|Y\|_{F}\big/\|\widetilde{Y}\|_{F}.$ (3)
This scaling encourages the reference embedding to match the current scale of the $t$-SNE embedding during optimization, making the global and local objectives more compatible. Note that $\alpha$ was treated as a constant in each gradient descent iteration. We discuss alternative scaling options in Appendix B.
The first term of Equation 2 enforces the preservation of local structure in the embedding by minimizing the $t$-SNE loss (Equation 1), while the regularization term ensures the preservation of global structure by encouraging the embedding to resemble the (scaled) PCA embedding $\widetilde{Y}$. This setup allows for local adjustments by the $t$-SNE loss, while the quadratic penalty prevents large deviations that would distort the global layout. The regularization strength $\lambda$ controls the impact of the PCA embedding on the final embedding, thereby enabling a trade-off between local and global structure preservation. For $\lambda=0$, DREAMS produces a standard $t$-SNE embedding since the loss term (Equation 2) reduces to the $t$-SNE objective. In the other limiting case of regularization strength $\lambda=1$, the objective becomes the regularizer without the $t$-SNE loss whose optimum is a (scaled) PCA embedding. For intermediate $\lambda\in(0,1)$ the objective is a weighted mean between the $t$-SNE objective and the regularizer. Empirically, we found $\lambda=0.15$ to be a suitable value for combining the individual strengths of global structure preservation of PCA with the local structure preservation of $t$-SNE, without substantially compromising either aspect.
This framework is not limited to using the PCA embedding as the global reference embedding $\widetilde{Y}$. Depending on the characteristics of the dataset, other methods with good global structure preservation, such as MDS, can also be used as an effective choice for $\widetilde{Y}$. We will refer to the version of DREAMS that uses the SQuadMDS embedding for regularization as DREAMS-MDS. By default, we use PCA instead of MDS because PCA is faster and is inherently interpretable due to its linearity.
Our implementation is based on the openTSNE library (Poličar et al., 2024), where we added the regularization term to the gradient-descent optimization. We used default openTSNE hyperparameters and initialized the embedding with the regularization embedding $\widetilde{Y}$, which was rescaled so that its first dimension had a standard deviation of $10^{-4}$ (as is default in openTSNE).